Derive the IM and IS from the following equation.
Commodity market y=c+I c=200+2/5y and I=1199-12r
Money market
MDT=1/2y
MDS=100-mr
MS=1500
1
Expert's answer
2021-07-28T13:35:02-0400

IS equation gives the good market equilibrium represented by:

Total Income=Aggregate Expenditure= Consumption +Investment.

"y=c+I"

"c=200+\frac{2}{5}y"

"I=1199-12r"

"y=(200+\frac{2}{5}y)+1199-12r"

"y=\frac{2}{5}y+1399-12r"

"y-\frac{2}{5}y=1399-12r"

"y(1-\frac{2}{5})=1399-12r"

"y=\frac{1399-12r}{\frac{3}{5}}"

"y=2332-20r" ...IS equation


LM equation gives the money market equilibrium represented by:

Money Demand =Money Supply

"(\frac{M}{P})^d=(\frac{M}{P})^s=d_1y-d_2r=\frac{M}{P}"

"r=(\frac{\frac{M}{P}}{d_2})+(\frac{d_1}{d_2})y" …LM equation